OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2

OpenStudy (anonymous):

(1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (anonymous):

(1, 5)

OpenStudy (anonymous):

thank you can you help me with another>]?

OpenStudy (anonymous):

@onaogh

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - <=5 y -x - <=4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (anonymous):

@onaogh

OpenStudy (anonymous):

(1, -10)

OpenStudy (anonymous):

thanks @onaogh i have 3 more

OpenStudy (anonymous):

The graph below represents which system of inequalities?

OpenStudy (anonymous):

y <=-2x + 3 y<= x + 3 y >=-2x + 3 y>= x + 3 y <=-3x + 2 y<= -x + 2 y > -2x + 3 y > x + 3

OpenStudy (anonymous):

y <=-2x + 3 y<= x + 3

OpenStudy (anonymous):

The graph below represents which system of inequalities?

OpenStudy (anonymous):

y > 2x - 3 y > -x - 3 y < 2x - 2 y < -x + 3 y <=2x - 2 y > -x + 3 None of the above

OpenStudy (anonymous):

y <=2x - 2 y > -x + 3

OpenStudy (jdoe0001):

how do you know that @onaogh ?

OpenStudy (anonymous):

from the graph

OpenStudy (anonymous):

Choose the graph below that represents the following system of inequalities. y >=-3x + 1 y >=x + 3

OpenStudy (jdoe0001):

the graph doesnt' say anything about 2 equations unless of course you draw them and get their equation

OpenStudy (anonymous):

@onaogh

OpenStudy (jdoe0001):

@essencedw are you planning to do your homework, or just have @onaogh do it for you? I think he's helped you enough so far that you can do it

OpenStudy (jdoe0001):

it helps you none you know

OpenStudy (anonymous):

flvs4.jpg

OpenStudy (anonymous):

thankyou so much :) @onaogh

OpenStudy (anonymous):

@jdoe0001 i didnt ask you i asked him/her so shhh and get out thanks

OpenStudy (anonymous):

jdoe0001 have a point, homeworks help you to better understand what you learn in school if you try to attempt solving them. solving homeworks yourself = better understanding = better grades